OpenStudy (anonymous):

find the equation of the locus of a point which moves so that its distance from the line 6x+2y=13 is always twice its distance from the line y=3x+8

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Do you know how to find the distance from a point to a line?

OpenStudy (anonymous):

use sq.rt(x^2+y^3)

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

hmm no not quite

jimthompson5910 (jim_thompson5910):

this page https://en.wikipedia.org/wiki/Distance_from_a_point_to_a_line offers a great simple formula for finding the distance https://upload.wikimedia.org/math/2/a/e/2ae89917912e4c37c8673d56ac84fa81.png

jimthompson5910 (jim_thompson5910):

keep in mind that they write the standard form equation as `ax+by+c = 0` so you'll have to turn `6x+2y=13` into `6x+2y-13=0`

OpenStudy (anonymous):

ok and 3x-y+8=0

jimthompson5910 (jim_thompson5910):

For some reason, the image in the second link is very small It should say this \[\Large \text{distance}(ax+by+c=0,(x_0,y_0)) = \frac{|ax_0+by_0+c|}{\sqrt{a^2+b^2}}\]

jimthompson5910 (jim_thompson5910):

Good, `y=3x+8` would turn into `3x-y+8 = 0`

jimthompson5910 (jim_thompson5910):

Let's make P be some unknown point (x,y)

jimthompson5910 (jim_thompson5910):

The distance from P to the line L1 `6x+2y-13=0` is \[\Large d(P,L1) = \frac{|ax_0+by_0+c|}{\sqrt{a^2+b^2}}\] \[\Large d(P,L1) = \frac{|6x+2y-13|}{\sqrt{6^2+2^2}}\] \[\Large d(P,L1) = \frac{|6x+2y-13|}{\sqrt{40}}\]

jimthompson5910 (jim_thompson5910):

The distance from P to the line L2 `3x-y+8 = 0` is \[\Large d(P,L2) = \frac{|ax_0+by_0+c|}{\sqrt{a^2+b^2}}\] \[\Large d(P,L2) = \frac{|3x-y+8|}{\sqrt{3^2+(-1)^2}}\] \[\Large d(P,L2) = \frac{|3x-y+8|}{\sqrt{10}}\]

jimthompson5910 (jim_thompson5910):

`find the equation of the locus of a point which moves so that its distance from the line 6x+2y=13 is always twice its distance from the line y=3x+8` This means \[\Large d(P,L1) = 2*d(P,L2)\]

jimthompson5910 (jim_thompson5910):

\[\Large d(P,L1) = 2*d(P,L2)\] \[\Large \Large \frac{|6x+2y-13|}{\sqrt{40}} = 2*\frac{|3x-y+8|}{\sqrt{10}}\] \[\Large \Large \left(\frac{|6x+2y-13|}{\sqrt{40}}\right)^2 = \left(2*\frac{|3x-y+8|}{\sqrt{10}}\right)^2\] \[\Large \Large \frac{(6x+2y-13)^2}{40}= 4*\frac{(3x-y+8)^2}{10}\]

OpenStudy (anonymous):

\[(6x+2y-13)^2=16(3x-y+8)^2\]

OpenStudy (anonymous):

\[(6x+2y-13)(6x+2y-13)=16(3x-y+8)(3x-y+8)\]\[36x^2+24xy+4y^2-156x-52y+169=16(9x^2-6xy+48x+y^2-16y+64)\]

OpenStudy (anonymous):

@jim_thompson5910